OpenStudy (anonymous):

write the equation of the line for m=5 and point (3,8)

OpenStudy (amistre64):

which equation?

OpenStudy (amistre64):

there are at least 4 forms of a line equation that I am aware of

OpenStudy (anonymous):

y=mx+b

OpenStudy (amistre64):

then: y = mx -mPx + Py y = 5x -5(3) + 8

OpenStudy (anonymous):

where p is?

OpenStudy (amistre64):

point (Px,Py) is the given point

OpenStudy (amistre64):

easier to type then all those y1 y0 x0 stuff

OpenStudy (amistre64):

y - Py = m(x-Px) y-Py = mx -mPx y = mx - mPx + Py

OpenStudy (anonymous):

true thanks